Adult tickets for a play cost $17 and child tickets cost $2. If there were 32 people at a performance and the theater collected $364 from ticket sales, how many adults and how many children attended the play?
Answer by fractalier(2101)

Let a = # of adult tickets. Their value would then be 17a.
Let c = # of child tickets. Their value would be 2c.
Now set up the two equations.
a + c = 32
17a + 2c = 364
Double the first equation and subtract it from the second...
17a + 2c = 364
2a + 2c = 64
and we get
15a = 300
a = 20 and then of course
c = 12